OpenStudy (anonymous):

Solve this system of equations using elimination x + 2y = 13 4x + 8y = -9

OpenStudy (anonymous):

is this algebra 1?

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

give me a minute..

OpenStudy (anonymous):

ok @GamerGirl.1

OpenStudy (anonymous):

First decide whether there to add or subtract the equations.Remember that the addition or subtraction should eliminate one variable. In the system above,the x in each equation is positive,so we need to subtract,that is,change all the signs of the terms in the second equation

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

?

OpenStudy (mathmale):

Solve this system of linear equations: x + 2y = 13 4x + 8y = -9 Multiply the first equation (every term!) by -4. You'll see why, in a moment. What do you get? Type your result here.

OpenStudy (anonymous):

-4x-8y=-52

OpenStudy (anonymous):

@mathmale I need to keep atleast one variable?

OpenStudy (mathmale):

Good. 4x + 8y = -9 (this is your second equation) -4x - 8y = -52 (this is your modified first equation) What would happen if you were now to add like terms together? (Add 4x-4, 8y-8y, etc.)

OpenStudy (anonymous):

So the answer is no solutions?

OpenStudy (solomonzelman):

yes, correct. No solution is the answer, because you can't have "4x+8y" be equal to -9 and at the same time be equal to 52